For pravinmandora

Assume that you just graduated from GW and are employed at an investment bank making $120,000 (after-tax) per year, and you expect to make the same amount for each of the next 5 years. A classmate from MAT62, who knows what a hard party person you were, gives you a call and tries to convince you to join forces with him on a project. The project is to produce a new type of vodka, “Hangover’s over”, that won’t make one feel hangover the next morning no matter how many bottles one drinks. If you decide to join, you will have to quit your current job, and work full-time on the project. The project requires an initial investment of $400,000 in production equipment, which can be depreciated straight-line over 5 years to a salvage value of $80,000. You own a house that you’re currently renting out for $24,000 a year and you’re planning to use the house as your office if you join the project. You expect to sell 20,000 bottles of the “Hangover’s over!” at $40 per unit each year for the next 5 years. The production cost is $15 per unit, and fixed costs associated with the project are estimated to amount $150,000 per year. Assume that your tax rate is 40% and your discount rate for projects with similar risk is 12%. Should you accept the project?
Answer #1

Opportunity cost per year = 120000 + 24000 = 144000

Total revenue per year = 20000*40 = 800000
Production cost = 20000*15 = 300000
Fixed cost = 150000
Depreciation = (400000-80000)/5 = 64000

Profit per year = 800000-300000-150000-64000 = 286000
Profit after tax = 286000*(1-40%) = 171600
Discounted cash flow for fist year = 171600/1.12 = 153214.3

Since the discounted cash flow is more than the total opportunity cost, the project should be accepted.